OpenStudy (anonymous):

Which ordered pair is a solution to the inequality y > x + 1? (1, 5) (0, -2) (5, -1) (2, 2)

OpenStudy (anonymous):

subtract x from both sides: y-x>1 therefore , from the choices given, that satisfy y-x>1 is the answer, got it?

OpenStudy (anonymous):

ummmm

OpenStudy (anonymous):

no :/ sorry

OpenStudy (anonymous):

Im just so bad at these :/

OpenStudy (anonymous):

5-1 =? -2-0=? -1-5=? 2-2=? of the above which solution is greater than 1?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

then that's your solution that satisfy's the inequality: y>x+1

OpenStudy (anonymous):

O ok thank you for explaining and working trough it with me :)

OpenStudy (anonymous):

you welcome